OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3

OpenStudy (anonymous):

x=1+4y..equ (3) substitute d value of x in equation 2 which is 2(1+4y)-9y=3 2+8y-9y=3 2-y=3 -y=3-2 -y=1 y=-1. then substitute the current value of y in equation 1 x-4y=1 x-4(-1)=1 x+4=1 x=1-4 x=-3 so x=-3,y=-1

OpenStudy (anonymous):

1st. You have to choose one of the two equations (I'll choose the first one) and express x in terms of y: x - 4y = 1 x = 1 + 4y 2nd. Substitute x in the second equation: 2x - 9y = 3 2 * (1 + 4y) - 9y = 3 2 + 2*4y - 9y = 3 2 + 8y - 9y = 3 2 - y = 3 2 - 3 = y y = -1 3rd. You know y, you can find x: x = 1 + 4y x = 1 + 4 * (-1) x = 1 - 4 x = -3